OpenStudy (anonymous):

will medal Which equation illustrates the multiplication property of equality?

OpenStudy (skullpatrol):

Do you know what this property is?

OpenStudy (anonymous):

A.-12=y times 3 -12 divided by 3=y times 3 divided by 3 -4=y B.-12=y-3 -12+3=y-3+3 -9=y C.-12=y divided by 3 -12 times 3=y divided by 3 times 3 -36=y D.-12=y+3 -12-3=y+3-3 -15=y

OpenStudy (anonymous):

@skullpatrol i think its c

OpenStudy (skullpatrol):

You are right! :D

OpenStudy (skullpatrol):

If a, b, and c are any real numbers and a = b, then ca = cb and ac = bc.

OpenStudy (skullpatrol):

If equal numbers are multiplied by the same number, the products are equal.